OpenStudy (anonymous):

John bought new jeans, but they were too long. Ratio length of jeans to Height of John is 5:8. Mum cut 4cm from jeans, by what the whole ratio was smaller by 4%. What is the height of John?

OpenStudy (anonymous):

okay u see the ratio changed by 4%

OpenStudy (anonymous):

the new ratio is X is 4% less than 5/8 so

OpenStudy (anonymous):

X=(1-0.04) * 5/8 solve for x

OpenStudy (anonymous):

then u can work out how big john must have been originally for 4cm cut of pants to bring this change

OpenStudy (anonymous):

Could you either help me with that? Just write the answer and how you got it, I will try to work it out with your help and then check if its same, okay? I wont cheat!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let D me the length of johns jeans at first

OpenStudy (anonymous):

no lets 5*D be the length

OpenStudy (anonymous):

5D/8D (5D-4)/(8D)=(1-0.04)*5/8

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

no, I cant work out the answer..